OpenStudy (anonymous):

Let f(x) = x + 8 and g(x) = x2 – 6x – 7. Find f(g(2)). –7 –3 10 33

OpenStudy (anonymous):

f(x) = x + 8 g(x) = x^2 - 6x - 7 f(g(2)) = (2^2 - 6(2) - 7) + 8 = 4 - 12 - 7 + 8 = -7